OpenStudy (anonymous):

Will fan and medal A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 30°, what is the area of the opening? 14.9in^2 16.0in^2 19.3in^2 20.8in^2

OpenStudy (anonymous):

@jim_thompson5910 can you help? Or do you know anyone who can help?

jimthompson5910 (jim_thompson5910):

\[\Large A(\theta) = 16*\sin(\theta)*(\cos(\theta) + 1)\] \[\Large A(30^{\circ}) = 16*\sin(30^{\circ})*(\cos(30^{\circ}) + 1)\] \[\Large A(30^{\circ}) = 16*\frac{1}{2}*(\frac{\sqrt{3}}{2} + 1)\] \[\Large A(30^{\circ}) \approx 16*0.5*(0.866025 + 1)\] \[\Large A(30^{\circ}) \approx ???\]

OpenStudy (anonymous):

So the answer would be A?

jimthompson5910 (jim_thompson5910):

yeah I'm getting 14.9282 which rounds to 14.9

OpenStudy (anonymous):

That's what I got :) Thank you so much